# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ Appendix I Ablation: Different learning rates for VI
Laplace Approximation 1e-2 1.000 ($\pm$ 0.000) 2.342 (±($\pm$( ± 0.390) 2.121 ($\pm$ 0.100) 1.000 ($\pm$ 0.000) 2.134 (±($\pm$( ± 0.107) 2.095 ($\pm$ 0.062)
Laplace Approximation 1e-3 1.000 ($\pm$ 0.000) 2.341 (±($\pm$( ± 0.389) 2.121 ($\pm$ 0.100) 1.000 ($\pm$ 0.000) 2.133 (±($\pm$( ± 0.108) 2.095 ($\pm$ 0.062)
Laplace Approximation 1e-4 1.000 ($\pm$ 0.000) 2.341 (±($\pm$( ± 0.389) 2.121 ($\pm$ 0.100) 1.000 ($\pm$ 0.000) 2.133 (±($\pm$( ± 0.108) 2.095 ($\pm$ 0.062)
VI: DiagonalNormal 1e-2 0.892 ($\pm$ 0.074) 0.921 (±($\pm$( ± 0.374) 1.411 ($\pm$ 0.174) 0.889 ($\pm$ 0.062) 0.819 (±($\pm$( ± 0.343) 1.339 ($\pm$ 0.190)
VI: DiagonalNormal 1e-3 0.966 ($\pm$ 0.024) 1.588 (±($\pm$( ± 0.540) 1.672 ($\pm$ 0.203) 0.981 ($\pm$ 0.017) 1.685 (±($\pm$( ± 0.331) 1.739 ($\pm$ 0.139)
VI: DiagonalNormal 1e-4 0.971 ($\pm$ 0.010) 1.572 (±($\pm$( ± 0.300) 1.666 ($\pm$ 0.081) 0.849 ($\pm$ 0.030) 0.575 (±($\pm$( ± 0.127) 1.221 ($\pm$ 0.098)
VI: MultivariateNormal 1e-2 0.725 ($\pm$ 0.064) 0.523 (±($\pm$( ± 0.242) 1.114 ($\pm$ 0.261) 0.625 ($\pm$ 0.051) 0.470 (±($\pm$( ± 0.066) 0.918 ($\pm$ 0.119)
VI: MultivariateNormal 1e-3 0.964 ($\pm$ 0.008) 1.455 (±($\pm$( ± 0.327) 1.617 ($\pm$ 0.100) 0.853 ($\pm$ 0.052) 0.634 (±($\pm$( ± 0.266) 1.238 ($\pm$ 0.151)
VI: MultivariateNormal 1e-4 0.984 ($\pm$ 0.005) 1.848 (±($\pm$( ± 0.324) 1.773 ($\pm$ 0.079) 0.899 ($\pm$ 0.020) 0.807 (±($\pm$( ± 0.094) 1.345 ($\pm$ 0.079)
VI: Structured Normal 1e-2 0.734 ($\pm$ 0.063) 0.541 (±($\pm$( ± 0.254) 1.119 ($\pm$ 0.264) 0.670 ($\pm$ 0.047) 0.467 (±($\pm$( ± 0.086) 1.060 ($\pm$ 0.130)
VI: Structured Normal 1e-3 0.882 ($\pm$ 0.042) 0.719 (±($\pm$( ± 0.315) 1.335 ($\pm$ 0.149) 0.776 ($\pm$ 0.045) 0.473 (±($\pm$( ± 0.081) 1.064 ($\pm$ 0.131)
VI: Structured Normal 1e-4 0.890 ($\pm$ 0.027) 0.710 (±($\pm$( ± 0.290) 1.347 ($\pm$ 0.138) 0.771 ($\pm$ 0.049) 0.468 (±($\pm$( ± 0.078) 1.062 ($\pm$ 0.128)
VI: IAF 1e-2 0.840 ($\pm$ 0.036) 0.502 (±($\pm$( ± 0.262) 1.272 ($\pm$ 0.170) 0.614 ($\pm$ 0.045) 0.455 (±($\pm$( ± 0.048) 0.957 ($\pm$ 0.105)
VI: IAF 1e-3 0.797 ($\pm$ 0.065) 0.485 (±($\pm$( ± 0.556) 1.169 ($\pm$ 0.313) 0.619 ($\pm$ 0.036) 0.469 (±($\pm$( ± 0.064) 0.989 ($\pm$ 0.124)
VI: IAF 1e-4 0.803 ($\pm$ 0.068) 0.475 (±($\pm$( ± 0.535) 1.162 ($\pm$ 0.291) 0.612 ($\pm$ 0.034) 0.457 (±($\pm$( ± 0.055) 0.977 ($\pm$ 0.113)
Model LR Synthetic Evaluation Real-World Evaluation
@@
Laplace Approximation 1e-2 1.000 ($\pm$ 0.000) 3.449 (±($\pm$( ± 0.821) 1.773 ($\pm$ 0.539) 1.000 ($\pm$ 0.000) 2.703 (±($\pm$( ± 0.312) 0.362 ($\pm$ 0.017)
Laplace Approximation 1e-3 1.000 ($\pm$ 0.000) 4.288 (±($\pm$( ± 0.853) 2.263 ($\pm$ 0.732) 1.000 ($\pm$ 0.000) 2.896 (±($\pm$( ± 0.238) 0.376 ($\pm$ 0.022)
Laplace Approximation 1e-4 1.000 ($\pm$ 0.000) 4.252 (±($\pm$( ± 0.611) 2.122 ($\pm$ 0.430) 1.000 ($\pm$ 0.000) 2.805 (±($\pm$( ± 0.181) 0.368 ($\pm$ 0.017)
VI: DiagonalNormal 1e-2 0.998 ($\pm$ 0.002) 2.880 (±($\pm$( ± 1.046) 1.457 ($\pm$ 0.559) 0.944 ($\pm$ 0.008) 1.022 (±($\pm$( ± 0.067) 0.230 ($\pm$ 0.010)
VI: DiagonalNormal 1e-3 0.998 ($\pm$ 0.002) 2.973 (±($\pm$( ± 0.834) 1.465 ($\pm$ 0.540) 0.941 ($\pm$ 0.006) 0.997 (±($\pm$( ± 0.056) 0.229 ($\pm$ 0.010)
VI: DiagonalNormal 1e-4 1.000 ($\pm$ 0.001) 3.416 (±($\pm$( ± 0.761) 1.602 ($\pm$ 0.437) 0.943 ($\pm$ 0.009) 0.997 (±($\pm$( ± 0.057) 0.229 ($\pm$ 0.010)
VI: MultivariateNormal 1e-2 0.993 ($\pm$ 0.007) 2.969 (±($\pm$( ± 1.089) 1.506 ($\pm$ 0.659) 0.929 ($\pm$ 0.007) 0.957 (±($\pm$( ± 0.048) 0.224 ($\pm$ 0.010)
VI: MultivariateNormal 1e-3 0.996 ($\pm$ 0.004) 3.140 (±($\pm$( ± 0.910) 1.570 ($\pm$ 0.625) 0.934 ($\pm$ 0.009) 0.971 (±($\pm$( ± 0.054) 0.225 ($\pm$ 0.010)
VI: MultivariateNormal 1e-4 0.997 ($\pm$ 0.007) 3.464 (±($\pm$( ± 0.791) 1.639 ($\pm$ 0.426) 0.934 ($\pm$ 0.005) 0.962 (±($\pm$( ± 0.049) 0.225 ($\pm$ 0.010)
VI: Structured Normal 1e-2 0.998 ($\pm$ 0.002) 3.005 (±($\pm$( ± 0.871) 1.481 ($\pm$ 0.504) 0.947 ($\pm$ 0.005) 1.003 (±($\pm$( ± 0.066) 0.230 ($\pm$ 0.009)
VI: Structured Normal 1e-3 0.999 ($\pm$ 0.001) 3.244 (±($\pm$( ± 0.665) 1.619 ($\pm$ 0.559) 0.948 ($\pm$ 0.007) 1.033 (±($\pm$( ± 0.078) 0.232 ($\pm$ 0.009)
VI: Structured Normal 1e-4 0.999 ($\pm$ 0.001) 3.119 (±($\pm$( ± 0.612) 1.487 ($\pm$ 0.400) 0.943 ($\pm$ 0.007) 0.998 (±($\pm$( ± 0.056) 0.229 ($\pm$ 0.010)
VI: IAF 1e-2 0.939 ($\pm$ 0.040) 2.836 (±($\pm$( ± 0.293) 1.247 ($\pm$ 0.297) 0.944 ($\pm$ 0.008) 1.518 (±($\pm$( ± 0.048) 1.332 ($\pm$ 0.027)
VI: IAF 1e-3 0.927 ($\pm$ 0.047) 2.758 (±($\pm$( ± 0.342) 1.195 ($\pm$ 0.331) 0.949 ($\pm$ 0.009) 1.560 (±($\pm$( ± 0.031) 1.392 ($\pm$ 0.024)
VI: IAF 1e-4 0.842 ($\pm$ 0.038) 2.862 (±($\pm$( ± 0.296) 1.281 ($\pm$ 0.292) 0.943 ($\pm$ 0.008) 1.493 (±($\pm$( ± 0.039) 1.302 ($\pm$ 0.039)
Model LR Synthetic Evaluation Real-World Evaluation

### v2

@@ I.3 Gaussian Mixture Models
ICL (ours) 0.765 ($\pm$ 0.123) 0.767 ($\pm$ 0.727) 0.585 ($\pm$ 0.301) 0.614 ($\pm$ 0.074) 0.175 ($\pm$ 0.219) 0.310 ($\pm$ 0.138)
ICL (ours) 0.839 ($\pm$ 0.072) 0.707 ($\pm$ 0.658) 1.111 ($\pm$ 0.300) 0.768 ($\pm$ 0.033) 0.143 ($\pm$ 0.089) 0.411 ($\pm$ 0.094)
ICL (ours) 0.611 ($\pm$ 0.070) 0.089 ($\pm$ 0.114) 0.423 ($\pm$ 0.348) 0.576 ($\pm$ 0.027) 0.037 ($\pm$ 0.026) 0.257 ($\pm$ 0.044)
ICL (ours) 0.753 ($\pm$ 0.049) 0.171 ($\pm$ 0.153) 0.631 ($\pm$ 0.294) 0.762 ($\pm$ 0.015) 0.105 ($\pm$ 0.046) 0.597 ($\pm$ 0.104)
ICL (ours) 0.621 ($\pm$ 0.063) 0.067 ($\pm$ 0.080) 0.299 ($\pm$ 0.195) 0.610 ($\pm$ 0.045) 0.046 ($\pm$ 0.020) 0.242 ($\pm$ 0.038)
ICL (ours) 0.532 ($\pm$ 0.019) 0.016 ($\pm$ 0.008) 0.590 ($\pm$ 0.066) 0.556 ($\pm$ 0.017) 0.035 ($\pm$ 0.015) 0.504 ($\pm$ 0.038)
For GLMs (Table 42), ICL achieves significantly better results, with notable improvements in C2ST. In Scenario 1, synthetic C2ST drops from 0.992 to 0.765 and real-world C2ST from 0.980 to 0.614. Similarly, Scenario 3 shows substantial gains, with synthetic C2ST improving from 0.997 to 0.611 and real-world C2ST from 0.983 to 0.576. These trends extend to metrics like $\mathcal{W}_{2}$, where ICL yields consistent reductions.
Scenario Model Synthetic Evaluation Real-World Evaluation
@@
ICL (ours) 0.552 ($\pm$ 0.028) 0.034 ($\pm$ 0.034) 0.289 ($\pm$ 0.083) 0.606 ($\pm$ 0.038) 0.068 ($\pm$ 0.069) 0.265 ($\pm$ 0.078)
ICL (ours) 0.542 ($\pm$ 0.006) 0.017 ($\pm$ 0.006) 0.244 ($\pm$ 0.033) 0.622 ($\pm$ 0.032) 0.098 ($\pm$ 0.039) 0.287 ($\pm$ 0.046)
ICL (ours) 0.537 ($\pm$ 0.023) 0.024 ($\pm$ 0.021) 0.259 ($\pm$ 0.088) 0.609 ($\pm$ 0.019) 0.124 ($\pm$ 0.037) 0.179 ($\pm$ 0.018)
ICL (ours) 0.684 ($\pm$ 0.060) 0.198 ($\pm$ 0.141) 0.918 ($\pm$ 0.246) 0.988 ($\pm$ 0.003) 1.764 ($\pm$ 0.026) 1.248 ($\pm$ 0.008)
ICL (ours) 0.535 ($\pm$ 0.016) 0.021 ($\pm$ 0.011) 0.279 ($\pm$ 0.060) 0.886 ($\pm$ 0.017) 1.207 ($\pm$ 0.101) 1.002 ($\pm$ 0.042)
ICL (ours) 0.543 ($\pm$ 0.021) 0.023 ($\pm$ 0.015) 0.345 ($\pm$ 0.173) 0.666 ($\pm$ 0.020) 0.200 ($\pm$ 0.034) 0.224 ($\pm$ 0.014)
Scenario Model Synthetic Evaluation Real-World Evaluation